the primeter of a rectangle is 36 cm and its area is 80 sq.cm. find its duimensions

Answer by mananth(16946)

perimeter = 36
2*(L+W)=36
/2
L+W=18
..
ARea =L*W
L*W = 80
L= 80/W
plug value of L in first equation
80/W + W = 36
(80+W^2)/W=36
80+W^2 = 36W
w^2-36W+80=0
use the quadratic formula to find the roots
x1= ((36+sqrt(976))/2
X1=33.6
..
X2=((36-sqrt(976))/2
X2=2.38
Dimensions are 33.6 by 2.4 cm
...
m.ananth@hotmail.ca